OpenStudy (angelwings996):

What is the number of real solutions? -11x^2 = x + 11

OpenStudy (dumbcow):

first set equal to zero second find discriminate (b^2 -4ac) third, if discriminate >0 : 2 real solutions if discriminate =0: 1 real solution if discriminate <0 : no real solutions

OpenStudy (angelwings996):

would the equation be -11x^2 - x - 11 = 0 ?

OpenStudy (dumbcow):

yes or you can get rid of the neg --> 11x^2 +x+11 = 0

OpenStudy (angelwings996):

Okay, now what would I do?

OpenStudy (dumbcow):

find Discriminant http://en.wikipedia.org/wiki/Discriminant

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then calculate \(b^2-4ac=...?\)

OpenStudy (angelwings996):

don't I already have a, b, and c so would I just substitute them into b^2 - 4ac ?

hartnn (hartnn):

what u got, a=..? b=...?c=..?

OpenStudy (angelwings996):

doesn't a = 11, b = 1, and c = 11 ?

hartnn (hartnn):

thats correct, now as @dumbcow said, u just need to find whether b^2-4ac is positive or negative...

OpenStudy (angelwings996):

so if I plug it in would it be like this .... \[1 ^{2} - 4(11)(11)\]

hartnn (hartnn):

yes, go on.

OpenStudy (angelwings996):

okay so \[1^{2} = 1 and -4(11) = -44(11) = -484\] Then you would subtract -484 from 1 which gives you -483, right ?

hartnn (hartnn):

yes, so how many real solutions ?

OpenStudy (angelwings996):

since -483 is less than 0 would it be no real solutions?

hartnn (hartnn):

absolutely correct :)

OpenStudy (angelwings996):

Okay thank you so much for both of your helps!

hartnn (hartnn):

welcome ^_^